# Explaining Explainability: Understanding Concept Activation Vectors

### v1

@@ 13.3 DeepDream
DeepDream [30] is a feature visualisation tool. It starts from an image, generated by sampling noise from a random uniform distribution and then iteratively updates the input image to maximise the L2 norm of activations of a particular layer. We use a similar approach but instead maximise the dot product between the activations and a CAV: ${\bm{v}}_{c,l}\cdot{\bm{a}}_{l}$. In Figures 18 and 19, we show these visualisations for a selection of ImageNet CAVs for successive layers in a ResNet-50.
@@ 14 Entanglement Experiment Details
In our visualisations (Fig. 4 and Sec. 14.1) we compare multiple CAVs for each concept, each with a different random probe dataset, denoted by $r$. This allows us to see how similar CAVs for the same concept are on repeat training runs. Each value in the visualisation is the mean cosine similarity between the concepts on its corresponding x and y axis labels between all CAVs which do not have the same random probe dataset, i.e.:
@@ 14.2 Polysemanticity
In Figures 4 and 22 we find that the vector representations of mutually exclusive concepts are anti-correlated with each other. Each concept vector does not just mean, for example, red. It means red, not green and not blue.
Elhage et al. [13] also find evidence for polysemantic representations. However, they found individual neurons which were polysemantic, whereas here vectors, i.e., groups of neurons, are polysemantic. In addition, the reasoning is different. The polysemanticity discussed in Elhage et al. [13] is caused by sparse features being compressed into fewer neurons than there are features. Here, we do not have sparse features and have more neurons than features. Instead, the polysemanticity is caused by associations between the concepts and the optimisation process favouring negatively correlated representations for mutually exclusive concepts.
@@ 14.3 Dot product distributions
The definition of entangled concepts in Eq. 6 uses the dot products of a CAV trained on concept $c_{1}$ with the the activations of a probe dataset for a different concept $c_{2}$. Figure Fig. 24 shows the distribution of dot products for a selection of CAVs, model training datasets and concept probe datasets. We use the same set of Elements datasets as in Sec. 6.2: $\mathbb{E}_{1}$, $\mathbb{E}_{2}$, $\mathbb{E}_{3}$, where $\mathbb{E}_{1}$ has no association between red and triangle and $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$ have successively stronger associations between the concepts. This changing association is apparent in the dot product distributions. For $\mathbb{E}_{1}$, the dot products of the test probe datasets differing from the CAV concept do not differ significantly from the dot products for random images (the negative probe dataset). Whereas, for $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$, the random distribution is shifted lower than for either CAV, even if the CAV is labelled differently from the test dataset. This shows that the red and triangle CAVs are entangled for these datasets/models.
@@ 15.6 Spatially Dependent TCAV Scores
Figure 33: Examples of spatially dependent TCAV scores in the spatially dependent version of Elements. Each subfigure is a separate class. The standard deviation is shown in black for significant results and red for insignificant results. The mean TCAV score for random CAVs are shown as horizontal black lines.
@@ 16 Further Related Work
Recent work highlighted problems with concept-based explanation methods. Ramaswamy et al. [34] showed that using different probe datasets to interpret the same model can lead to different explanations for the same concept. Similarly, Soni et al. [38] showed these methods to be sensitive to the random seed used to sample images for the negative set. Our work complements this research by investigating the underlying properties of concept vectors and how they may cause problems when interpreting concept-based explanations.
Extensions to the original TCAV have been suggested, attempting to improve aspects of the original method. For instance, Ghorbani et al. [17] automate concept discovery by using super-pixels and clustering, removing the need to handcraft a probe dataset. Zhang et al. [48] and Schrouff et al. [37] change how CAVs are created to produce local and global explanations. However, these methods still use vectors to represent concepts [3]. As such, our work is still applicable to each of the extensions.
Table 3: Examples in computer vision and medical imaging research, where consistency, entanglement and spatial dependence may impact analyses. We use the following abbreviations: skin cancer (SC), skin lesions (SL), breast cancer (BC), histology (H) CIFAR-10/100 (CF) [25], COCO (CO) [27], CUB (CB) [40], Places365 (Pl) [49], Waterbirds (WB) [36], ImageNet (Im) [12]
@@
Consistency [42, 34, 16, 45, 18, 28]
Entanglement [42, 34, 16, 45, 18, 19, 28, 33]
Spatial Dependence [42, 34, 16, 45, 18, 28, 33]
17 Example Use-case
In §7 we provided recommendations to better use CAVs in practice. To provide a concrete example, we next examine the use-case of Yan et al. [42] which uses CAVs in the context of skin cancer diagnosis. In the next few sections we will demonstrate with this concrete example three recommendations, and show how these recommendations could have impacted the conclusions drawn.
Consistency
Entanglement
Spatial Dependence
17.1 Experiment Setup
As the dataset used by Yan et al. [42] is not publicly available, we reproduce their results using a similar publicly available dataset.
Model
We finetune a ResNet50 on the ISIC 2019 dataset [39, 8, 10] for the binary classification of melanoma. We use a binary cross entropy loss and the Adam optimiser [24], training until convergence of validation loss to achieve an area under the receiver operating characteristic curve (AUC) of 0.91 on the validation split.
CAVs
For the CAVs, as in Yan et al. [42], we use the derm7pt dataset [22]. There are $12$ clinical concepts which have been expertly labelled within the dataset. We hand labelled three additional concepts, which were used in [42], of dark corners, dark borders and ruler which are possible confounders for the model. We defined dark corners as any image with a circular aperture which left the corners of the image black, dark borders as any image containing rectangles of blacked out areas and ruler as any image containing a ruler. For each of the medical concepts, there are three labels: typical/regular, atypical/irregular, and absent. When training CAVs for these concepts, we used images with the label of ‘absent’ as the negative probe dataset. This was done because using random images as the negative set gave low-quality CAVs, with accuracies of $50$-$65$%. It is unclear from [42] what they used for negative sets. For training the CAVs we used 70 images per concept and used 30 different random seeds for the random negative probe set to get 30 CAVs per concept. Yan et al. [42] do not use TCAV, so this setup is different from the original paper.
17.2 Results
We first show that we can produce concept-based explanations for melanoma classification as in [42]. We then show how the use of our analysis tools impact the conclusions that can be drawn.
In Fig. 35, we obtain a high accuracy for the CAVs trained on confounders between $80$-$90\%$, but only $60$-$75\%$ for the medical concepts. We hypothesise that this is due to the simplicity of the confounding concepts. This is supported by the accuracy reported in [42] where they also obtained a lower CAV accuracy for the medical concepts. The accuracy for the confounding concepts of dark borders and dark corners drops in later layers. This is likely due to the spatial nature of the concepts – an idea further supported by Fig. 38 where the CAVs in later layers have reduced spatial distinction.
Figure 35: Mean test accuracy for the linear classifiers from which the CAVs are generated for all concepts analysed on the melanoma datasets (split by concept type - confounders, regular, irregular)
Consistency – Consistent TCAV scores across layers indicate reliable explanations. The TCAV scores in Fig. 36 are consistent across layers, providing more confidence in using them to explain the model’s behaviour. In terms of understanding the model, the scores provide some evidence that it operates similar to human experts, as the TCAV scores for the atypical/irregular medical concepts are high for the malignant class, as expected, and the confounding concepts are often not significant, providing no evidence that the model is sensitive to the confounders.
Figure 36: TCAV scores for the melanoma use-case. Top: medical concepts, Bottom: potential confounders. The standard deviation is shown in black for significant results and red for insignificant results. The mean TCAV score for random CAVs are shown as horizontal black lines.
Entanglement – Unexpected similarities between concepts suggest further experiments need to be performed. Interestingly, in Fig. 37, the CAVs with opposed meanings often appear to be the most similar to each other. For example, regular vascular structures has a negative or zero similarity with all concepts except irregular vascular structures (with a similarity of $0.11$). This could be due to our choice in negative set for the probe datasets as these two concepts share the same negative set. If we were to continue analysing these results, this hypothesis should be explored and alternative approaches attempted. Equally, we would require a domain expert to compare which concepts are most similar to each other to check whether the relative similarities of the concepts make sense.
Figure 37: Mean pairwise cosine similarities between 30 CAVs for different concepts from the melanoma use-case.
Spatial Dependence – The spatially dependent confounders are better represented in earlier layers. The spatial norms and means in Figures 38 and 39 show a clear spatial dependence in the center for each of the medical concepts across all layers. This aligns with expectations, as the dataset requires the skin lesion to be centered in the image and each concept is related to the appearance of the lesion. The dark corners and dark borders concepts, however, show deviations to this pattern in the earlier layers, with dark corners having high spatial norms in the corners and dark borders high spatial norms in the top. This is desirable, as these confounding concepts depend on features at the edge of the image, away from the lesion. Along with the accuracies in Fig. 35, this suggests that the CAVs in earlier layers better represent these spatially dependent concepts. For dark borders, we may want to examine the probe dataset and collect more images with dark borders in locations other than the top, as the CAV seems focused on the top of the feature space – this is even more apparent in the spatial means of Fig. 39.
Figure 38: Mean CAV spatial norms across 30 CAVs for a selection of concepts and layers in the melanoma use-case.
Figure 39: Mean CAV spatial means across 30 CAVs for a selection of concepts and layers in the melanoma use-case.

### v2

@@ 12.3 DeepDream
DeepDream (Mordvintsev et al., 2015) is a feature visualisation tool. It starts from an image, generated by sampling noise from a random uniform distribution and then iteratively updates the input image to maximise the L2 norm of activations of a particular layer. We use a similar approach but instead maximise the dot product between the activations and a CAV: ${\bm{v}}_{c,l}\cdot{\bm{a}}_{l}$. In Figures 22 and 23, we show these visualisations for a selection of ImageNet CAVs for successive layers in a ResNet-50.
@@ 12.4 Inconsistent TCAV Scores
12.5 TCAV Layer Consistency Score
At the end of § 6.1 we defined the TCAV layer consistency score as a measure of how often significant TCAV scores disagree on the direction of sensitivity for a specific concept and class across a set of layers. Here, we provide some example results for both the ImageNet and Elements datasets. In each case we train $30$ CAVs for each concept/layer pair to calculate the TCAV score across. For Elements, when analysing all 69 classes from the simple version of the dataset, we obtain a mean TCAV consistency score of $0.841$ across all concepts and all layers of the simple NN. $38/690$ sets of TCAV scores had a TCAV layer consistency score of less than $0.2$ and $158/690$ less than $0.5$. For ImageNet, we obtain a mean TCAV consistency score of $0.868$ across the last $8$ convolutional layers (from layer $3.0$ to layer $4.1$) for a range of concepts (striped, polka, meshed, road, sidewalk, car, bicycle, boat, person, anemone, water, waterfall, sea, sky, snow, tree) and classes (zebra, leopard, tiger, car wheel, acoustic guitar, academic gown). $7/96$ sets of TCAV scores had a TCAV layer consistency score of less than $0.2$ and $13/96$ had a score less than $0.5$. See Figure 26 for the full distributions.
Figure 26: Distribution of TCAV layer consistency score for the simple Elements dataset (left) and a selection of classes/concepts for a ResNet-50 trained on ImageNet (right).
@@ 13 Entanglement Experiment Details
In our visualisations (fig. 4 and section 13.1) we compare multiple CAVs for each concept, each with a different random probe dataset, denoted by $r$. This allows us to see how similar CAVs for the same concept are on repeat training runs. Each value in the visualisation is the mean cosine similarity between the concepts on its corresponding x and y axis labels between all CAVs which do not have the same random probe dataset, i.e.:
@@ 13.2 Polysemanticity
In Figures 4 and 27 we find that the vector representations of mutually exclusive concepts are anti-correlated with each other. Each concept vector does not just mean, for example, red. It means red, not green and not blue.
Elhage et al. (2022) also find evidence for polysemantic representations. However, they found individual neurons which were polysemantic, whereas here vectors, i.e., groups of neurons, are polysemantic. In addition, the reasoning is different. The polysemanticity discussed in Elhage et al. (2022) is caused by sparse features being compressed into fewer neurons than there are features. Here, we do not have sparse features and have more neurons than features. Instead, the polysemanticity is caused by associations between the concepts and the optimisation process favouring negatively correlated representations for mutually exclusive concepts.
@@ 13.3 Dot product distributions
The definition of entangled concepts in eq. 6 uses the dot products of a CAV trained on concept $c_{1}$ with the the activations of a probe dataset for a different concept $c_{2}$. Figure fig. 29 shows the distribution of dot products for a selection of CAVs, model training datasets and concept probe datasets. We use the same set of Elements datasets as in section 6.2: $\mathbb{E}_{1}$, $\mathbb{E}_{2}$, $\mathbb{E}_{3}$, where $\mathbb{E}_{1}$ has no association between red and triangle and $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$ have successively stronger associations between the concepts. This changing association is apparent in the dot product distributions. For $\mathbb{E}_{1}$, the dot products of the test probe datasets differing from the CAV concept do not differ significantly from the dot products for random images (the negative probe dataset). Whereas, for $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$, the random distribution is shifted lower than for either CAV, even if the CAV is labelled differently from the test dataset. This shows that the red and triangle CAVs are entangled for these datasets/models.
@@ 14.6 Spatially Dependent TCAV Scores
Figure 38: Examples of spatially dependent TCAV scores in the spatially dependent version of Elements for a ResNet-50. Each subfigure is a separate class. The standard deviation is shown in black for significant results and red for insignificant results. The mean TCAV score for random CAVs are shown as horizontal black lines.
14.7 Spatially Dependent CAVs in ViTs
Even though we cannot use spatial norms to visualise the spatial dependence of CAVs in a transformer based architecture, in this section we have some preliminary results suggesting that that spatially dependent CAVs can still be created for a ViT-B16 (Dosovitskiy et al., 2021). As for the ResNet-50, we test the model on the spatially dependent Elements dataset and found TCAV scores that vary by location.
We finetuned a ViT-B16 model pretrained on ImageNet for $50$ epochs on the spatially dependent version of Elements (i.e. there are some classes which depend on the location of the objects as well as which concepts are present). We used an exponentially decaying learning rate with initial learning rate of $0.0001$ and a $\gamma$ of $0.95$. Similarly to the previous section, we demonstrate that we obtain different TCAV scores for the left versions of a concept than the right, although there is substantially less variation in TCAV score than for the ResNet-50 (fig. 38). Future work should perform more extensive experiments to determine how the properties analysed in this paper are affected by architecture.
Figure 39: Spatially dependent TCAV scores in the spatially dependent version of Elements for a ViT-B16. The TCAV scores for the classes “striped triangles on the left” (top) and “striped triangles on the right” (bottom) are shown. The standard deviation is shown in black for significant results and red for insignificant results. The mean TCAV score for random CAVs are shown as horizontal black lines.
@@ 15 Further Related Work
Recent work highlighted problems with concept-based explanation methods. Ramaswamy et al. (2022a) showed that using different probe datasets to interpret the same model can lead to different explanations for the same concept. Similarly, Soni et al. (2020) showed these methods to be sensitive to the random seed used to sample images for the negative set. Our work complements this research by investigating the underlying properties of concept vectors and how they may cause problems when interpreting concept-based explanations.
Extensions to the original TCAV have been suggested, attempting to improve aspects of the original method. For instance, Ghorbani et al. (2019) automate concept discovery by using super-pixels and clustering, removing the need to handcraft a probe dataset. Zhang et al. (2020) and Schrouff et al. (2021) change how CAVs are created to produce local and global explanations. However, these methods still use vectors to represent concepts (Alain & Bengio, 2017). As such, our work is still applicable to each of the extensions.
The properties analysed in this paper are generally applicable. To give insight into when the various properties may be relevant, we performed a review of papers which use CAVs in medical imaging and computer vision research. In each case, we checked if the authors had done any checks related to layer consistency, entanglement or spatial dependence of CAVs. Almost no papers evaluated the effect of these properties on their results, and when they did it was by creating CAVs in multiple layers, providing some robustness to layer inconsistency. In Table 3, we provide a list of these papers, detailing any use-cases where our recommendations could have helped check the impact of CAV properties on results.
Table 3: Examples in computer vision and medical imaging research, where consistency, entanglement and spatial dependence may impact analyses. We use the following abbreviations: skin cancer (SC), skin lesions (SL), breast cancer (BC), histology (H) CIFAR-10/100 (CF) (Krizhevsky, 2009), COCO (CO) (Lin et al., 2014), CUB (CB) (Wah et al., 2011), Places365 (Pl) (Zhou et al., 2017), Waterbirds (WB) (Sagawa et al., 2020), ImageNet (Im) (Deng et al., 2009)
@@
Consistency Yan et al. (2023); Ramaswamy et al. (2022a); Fürböck et al. (2022); Ghosh et al. (2023); Lucieri et al. (2020); Yuksekgonul et al. (2023)
Entanglement Yan et al. (2023); Ramaswamy et al. (2022a); Fürböck et al. (2022); Ghosh et al. (2023); Graziani et al. (2020); Lucieri et al. (2020); Pfau et al. (2020); Yuksekgonul et al. (2023)
Spatial Dependence Yan et al. (2023); Ramaswamy et al. (2022a); Fürböck et al. (2022); Ghosh et al. (2023); Lucieri et al. (2020); Pfau et al. (2020); Yuksekgonul et al. (2023)